Let f(x) = |2x − 10| + 2 . a) Show that f is continuous at x = 5. [8] b) Show that f is not differentiable at x = 5
1
Expert's answer
2021-10-26T03:21:16-0400
"f(x) = |2x-10| + 2"

a)

"\lim\limits_{x\to5^-}f(x)=\lim\limits_{x\to5^-}(|2x-10| + 2 )"

"=\lim\limits_{x\to5^-}(10-2x + 2 )=12-2(5)=2"

"\lim\limits_{x\to5^+}f(x)=\lim\limits_{x\to5^+}(|2x-10| + 2 )"

"=\lim\limits_{x\to5^+}(2x-10 + 2 )=2(5)-8=2"

Then


"\lim\limits_{x\to5^-}f(x)=2=\lim\limits_{x\to5^+}f(x)=>\lim\limits_{x\to5}f(x)=2"

"f(5)=|2(5)-10|+2=0+2=2"

We have


"\lim\limits_{x\to5}f(x)=2=f(5)"

Therefore the function "f(x)" is continuous at "x=5."


b)


"f(5)=|2(5)-10|+2=0+2=2""\lim\limits_{h\to0^-}\dfrac{f(5+h)-f(5)}{h}=\lim\limits_{h\to0^-}\dfrac{|2(5+h)-10|+2-2}{h}"

"=\lim\limits_{h\to0^-}\dfrac{2|h|}{h}=\lim\limits_{h\to0^-}\dfrac{-2h}{h}=-2"




"\lim\limits_{h\to0^+}\dfrac{f(5+h)-f(5)}{h}=\lim\limits_{h\to0^+}\dfrac{|2(5+h)-10|+2-2}{h}"

"=\lim\limits_{h\to0^+}\dfrac{2|h|}{h}=\lim\limits_{h\to0^+}\dfrac{2h}{h}=2"

Since

"\lim\limits_{h\to0^-}\dfrac{f(5+h)-f(5)}{h}=-2"

"\not=2=\lim\limits_{h\to0^+}\dfrac{f(5+h)-f(5)}{h},"

then "\lim\limits_{h\to0}\dfrac{f(5+h)-f(5)}{h}" does not exist.

Therefore the function "f(x)" is not differentiable at "x=5."